Can someone please show me how to do this:
(s+3)(s-3)

Answer by benni1013(206)

Use the foil method:
(s+3)(s-3)
First:s*s=s^2
Outer:3*s=3s
Inner:-3*s=-3s
Last:3*-3=-9
s^2+3s-3s-9==>Simplify==>s^2-9